A mass is hung at the end of the spring length of 14 cm. The mass was pull down




and extend to the quarter of the original length of the springs and released so that




it oscillates vertically. If the mass took 0.125 s to go up and down from




equilibrium point:




a) The acceleration when the displacement is at maximum.

1
Expert's answer
2022-02-22T08:52:28-0500

Given:

"l_0=0.14\:\rm m"

"A=1/4l_0=0.035\:\rm m"

"T=0.125\:\rm s"


The angular frequency of oscillations

"\omega=\frac{2\pi}{T}=\frac{2\pi}{0.125}=50.24\;\rm rad/s"

The maximum acceleration

"a_{\max}=A\omega^2\\
=0.035*50.24^2=88\:\rm m/s^2"